OpenStudy (anonymous):

Let f(x) =2x^2. Simplify the expression f(x+h)-f(x)/h

Nnesha (nnesha):

replace x by (x+h) in f(x) equation \[\huge\rm f(x) = 2x^2\] replace x by (x+h)

OpenStudy (zehanz):

Because \(f(x)=2x^2\), you can start with: \(\dfrac{f(x+h)-f(x)}{h}=\dfrac{2(x+h)^2-2x^2}{h}\). Now look at this thing and think of what you can do with it...

OpenStudy (anonymous):

Is this going to be the answer 2(h+2x)?? @Nnesha